Velma is an Account Manager for Sherwin Williams paint. For the month, Velma sold paint that totaled $41,943 in sales. The cost of goods sold of this paint was 80%, so she generated 20% of sales as profit (or gross margin) for the company. Velma is paid 42% straight commission based on her profit. What is the total of her monthly compensation? (round to nearest dollar)
Answer #1

Solution:
Given in the Question
Total Sales = $41943
Cost of Good sold = 80% of the Sales i.e. ($41943*80%) = $33554.4
Profit on Sale = 20% of the Sales i.e. ($41943*20%) = $8388.6

Velma monthly compensation = 42% of Profit on Sale i.e. ($8388.6*42%) = $3523.212 or 3523
So Total monthly compensation = $3523